OpenStudy (anonymous):

can u help me solve 3(s+4) = 4s + 3(s-1)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (angela210793):

3s+12=4s+3s-3 12=4s-3 4s=15 s=15/4

OpenStudy (anonymous):

its for solving equations with variables on each side

OpenStudy (angela210793):

3s+12-4s-3s+3=0 15-4s=0 15=4s s=15/4